# Tadashi: Enabling AI-Based Automated Code Generation With Guaranteed Correctness

### v1

@@ 3.2.5. Auto-Tuning
For practical reasons, ML approaches used in auto-tuning (Mustafa, 2022; Ashouri et al., 2018) mostly rely on local search methods with focused search around particular regions in the search space, after aggressively pruning the search space. Similar to the exploitation in evolutionary algorithms, Tadashi can be setup to narrow down the search around the region for auto-tuning, based on the post-pruning range of interest.
@@ 3.2.6. ML-driven deployment of applications on new hardware
In addition to supporting the usage patterns of the ML approaches listed above, Tadashi can drive an ML approach to learnhardware specific optimisations on new hardware. After the training phase, applications ported to the new hardware can do inference on the ML model, leveraging Tadashi, to optimise the applications for the new hardware.
3.3. Extensibility of Tadashi
Tadashi is designed to be extensible w.r.t. the supported transformations, programming languages, and target hardware.
We implemented the set of transformations listed in Table 1. To achieve a larger coverage, further transformations (e.g. loop fission) could be added. This requires implementing a function which takes a schedule node as a parameter, performs the desired transformation and returns the transformed node (for details see Section 4.4.2). There are also features of ISL we did not explore yet, such as, expansion and extension nodes. A promising direction is to generate arbitrary partial schedules (which can be virtually any quasi-affine map444Namely isl_multi_union_pw_aff!.) for the band nodes in the schedule tree, providing access to space of all polyhedral transformations.
In essence, Tadashi’s core capabilities involve operating on ISL schedules, and it is not of importance what tool was used to generate them (the frontend) or how the output code is generated (the code generation backend). This means that while we currently use PET and ISL for those stages, which enables us to target C code, in the future we can integrate with compiler solutions such as Polly (Grosser et al., 2012) and Polygeist (Moses et al., 2021) to tap into their respective LLVM and MLIR ecosystems, or other tools which use ISL schedules as an intermediate representation (Vasilache et al., 2018; Verdoolaege et al., 2013) to gain additional capabilities w.r.t input languages or target hardware.
One potential scalability issue in utilising Tadashi for generating a large amount of runtime measurements is the availability of a wide range of inputs to run the generated code with. Benchmarks usually contain a small set of inputs for running the application, however, needing to run the entire application in order to time only specific small portions of it (scops) is wasteful. Since we already isolate these points of interest (scops), existing frameworks for automatic input generation (Ivanov et al., 2024; Kind et al., 2022) can be utilised to synthesise inputs which can be used to execute and time only the code in the scop.
@@ 4.1. The App Class
The App class (Listing 2 line 2) is completely written in Python and provided mostly for convenience. It’s main functions are to provide the source_path and include_paths properties, the former being the C source file which will be transformed, and the later a list of paths used to set the required environment variable,555Namely C_INCLUDE_PATH!. which is necessary, because when PET runs a compiler pass on the input file, it will throw an error if it cannot find an included header file.
@@ 4.2. SCoPs
The creation of the Scops object (line 3) sets things in motion in the C/C++ back-end of Tadashi. The input file is passed through PET’s C source transformations function666Namely pet_transform_C_source!. which detects all the SCoPs in the file, but instead of doing any transformations, the pet_scop objects passed to the callback are used to populate the global SCOP_INFO vector. Each entry in this vector is a scop_info_t struct, and stores the pointer to the pet_scop, the dependencies computed by ISL according to the instruction found in (Verdoolaege et al., [n. d.]) and using the stricter “may” dependencies,777Using isl_union_flow_get_may_dependence!. and pointer to the “current node”888Of type isl_schedule_node! in the schedule tree.
@@ 4.3. Schedule Tree and Navigation
Each transformation in Tadashi is defined on a node of a schedule tree, so line 4 in Listing 2 involves the following mechanism. First, scops[i] returns a Python Scop object which stores the index of the corresponding in SCOP_INFO entry on the C++ side. This Scop object wraps all the C++ functions exposed to Python by calling them with a scop_idx parameter from an object method with same name. These functions include navigation operations999Namely goto_root()!, goto_parent()!, and goto_child(idx)!. to update the current node to point to its parent, its child, or to the root of the schedule tree (for the selected scop_info_t object). The schedule_tree property on line 4 is dynamically generated, and exposes the schedule tree as a list of nodes by doing a depth-first traversal. During the traversal, we build a list of child indexes which enables Tadashi to locate any node by calling goto_child for each element in the list (implemented as the locate method of a Node). In addition to the navigation methods, Tadashi exposes wrappers to query properties of nodes, (such as node type, number of children, partial schedule, parameters, enclosing iterator variables etc.), and so enables the Python Node object to emulate the ISL schedule tree and its nodes in C++.
@@ 4.4.1. Implementation Details of Individual Transformations
To merge all band node under a sequence (FULL_FUSE), we take their partial schedules, intersect them with the corresponding filters and take the union. Then we introduce a new band node on top of the sequence101010Using isl_schedule_insert_partial_schedule!. and delete the original band nodes.
The 6 different SHIFT transformations, while simple on the surface, involve a tedious process involving the unpacking of the partial schedule, iterating through the sets of its domain and (re)constructing the adequate mupa111111Short for isl_multi_union_pw_aff!. object.
@@
4.4.2. Extending Tadashi With New Transformations
The entry point of a transformation is a function with a SCoP index121212By convention size_t scop_idx!. as its first argument (potentially followed by other arguments) in ctadashi.cc. Unless the transformation is trivial, the main implementation is added to transformations.c as a function which takes a schedule node (with the rest of the arguments) and returns the transformed schedule node. This function is then called from the entry-point function with the current node of the SCoP specified by the SCoP index. The call needs to be prefixed a pre- and followed by a post-transform (convenience) functions, which take care of some bookkeeping to be able to rollback properly. On the Python side, a class derived from TranfromInfo stores the name of the entry-point and implements methods which check and provide information for the availability of the transformation on a given node. Finally the TrEnum and global TRANSFORMATION dictionary connect the TransformInfo into the mechanism implemented by the Node’s transform method.
@@ 4.5. Legality Checking
After a transformation concludes, the legality check is invoked on the same node. The legality check implements the principles described in Section 2.2. Each step of the legality check in Algorithm 1 can be implemented with a single ISL function. Algorithm 1 is applied after every transformation with the exception of SET_PARALLEL. For the legality of parallel loops the entry in corresponding to the loop should be , and this can be achieved with a slight modification to Algorithm 1: we use the partial schedule instead of the complete schedule in steps 1 and 2, and in step 4 we compute and and check if they are both empty in step 5. Figure 4 shows two examples of the process of checking legality.
isl_union_map) from scop_info_t
scop_info_t
@@ 4.6. Code Generation
Apart from the obvious tweaking to use the modified schedule instead of the original one, and to make it cooperate with the bookkeeping setup we use in Tadashi’s back-end, the SET_PARALLEL transformation required additional modifications to the core code generation. Parallel loops are implemented as OpenMP parallel for loops, so the relevant pragma needs to be inserted. This is achieved by properly setting up the isl_ast_build and isl_ast_print_options objects, i.e., specifying custom after_each_mark and print_for callbacks for the AST build and print options object respectively.
@@ 5. Evaluation
We evaluate Tadashi using Random-RL, a mock ML model, verify the implementation and measure performace. All the measurements are done on a regular desktop with a 12 core AMD Ryzen 5 5600X 6-Core Processor and 64GiB of memory, running Arch Linux, with python 3.12.5-1, libisl 0.27-1, llvm 18.1.8-4 system packages, and Tadashi’s CMake script using the current master branch131313Git hash 824ab35b60bab089fb699e521e2e3bc6c07ce95a!. of PET141414From https://repo.or.cz/pet.git. tagged as pet 0.11.8.
5.1. Random-RL
To demonstrate the ease of use and performance of Tadashi, we show an example case of how Tadashi can be used in ML in an end-to-end fashion. We implement a mock RL script, Random-RL, described in Listings 3, with the agent randomly picking actions, which we evaluate on the Polybench (Reisinger, 2024) benchmark suit. Choosing a random action, consists of selecting a node in the schedule tree, a transformation, and parameters for that transformation. Technically, a random SCoP should also be chosen, but for simplicity, we assume that there is only one SCoP, which is conveniently true for the Polybench apps. The Random-RL agent keeps track of the last chosen node in the schedule tree, and the new node is selected by choosing randomly from its parent or one of it’s children (initially starting from the root). This is followed by a random sampling of all implemented transformations. The transformation might not be available for the selected node, so this process of picking random transformations and schedule nodes is repeated until a valid transformation-node pair is found. Finally, a valid set of transformation arguments is randomly picked, as described in Section 3.1.1.
5.2. Verification
@@
With the Random-RL agent successfully sampling the transformations, we verify the correct implementation of Tadashi. First, all Polybench benchmarks are compiled and executed with the small dataset151515Using -DSMALL_DATASET!. and with the output saved.161616Using -DPOLYBENCH_DUMP_ARRAYS!. In an entirely automated manner, we generate 3 random transformations using the Random-RL’s agent (as described above), rollback the illegal ones, compile, run, and compare the output of the transformed program to the original. Post-inspect confirmed the outputs to be identical.
5.3. Performance
@@ Acknowledgements.
- Ashouri et al. (2018) Amir H. Ashouri, William Killian, John Cavazos, Gianluca Palermo, and Cristina Silvano. 2018. A Survey on Compiler Autotuning using Machine Learning. ACM Comput. Surv. 51, 5, Article 96 (sep 2018), 42 pages. https://doi.org/10.1145/3197978
@@
- Cai et al. (2023) Zonghui Cai, Xiao Yang, MengChu Zhou, Zhi-Hui Zhan, and Shangce Gao. 2023. Toward explicit control between exploration and exploitation in evolutionary algorithms: A case study of differential evolution. Information Sciences 649 (2023), 119656. https://doi.org/10.1016/j.ins.2023.119656
@@
- Coello (2022) Carlos A. Coello Coello. 2022. Constraint-handling techniques used with evolutionary algorithms. In Proceedings of the Genetic and Evolutionary Computation Conference Companion (Boston, Massachusetts) (GECCO ’22). Association for Computing Machinery, New York, NY, USA, 1310–1333. https://doi.org/10.1145/3520304.3533640
- Cotroneo et al. (2024) Domenico Cotroneo, Alessio Foggia, Cristina Improta, Pietro Liguori, and Roberto Natella. 2024. Automating the correctness assessment of AI-generated code for security contexts. Journal of Systems and Software 216 (2024), 112113. https://doi.org/10.1016/j.jss.2024.112113
- Cummins et al. (2020) Chris Cummins, Zacharias V. Fisches, Tal Ben-Nun, Torsten Hoefler, and Hugh Leather. 2020. ProGraML: Graph-based Deep Learning for Program Optimization and Analysis. (March 2020). https://doi.org/10.48550/arXiv.2003.10536
@@
- Cummins et al. (2024) Chris Cummins, Volker Seeker, Dejan Grubisic, Baptiste Roziere, Jonas Gehring, Gabriel Synnaeve, and Hugh Leather. 2024. Meta Large Language Model Compiler: Foundation Models of Compiler Optimization. https://doi.org/10.48550/arXiv.2407.02524 arXiv:2407.02524 [cs]
- Dehaerne et al. (2022a) Enrique Dehaerne, Bappaditya Dey, Sandip Halder, Stefan De Gendt, and Wannes Meert. 2022a. Code Generation Using Machine Learning: A Systematic Review. IEEE Access 10 (2022), 82434–82455. https://doi.org/10.1109/ACCESS.2022.3196347
- Dehaerne et al. (2022b) Enrique Dehaerne, Bappaditya Dey, Sandip Halder, Stefan De Gendt, and Wannes Meert. 2022b. Code Generation Using Machine Learning: A Systematic Review. IEEE Access 10 (2022), 82434–82455. https://doi.org/10.1109/ACCESS.2022.3196347
- Du et al. (2024) Xueying Du, Mingwei Liu, Kaixin Wang, Hanlin Wang, Junwei Liu, Yixuan Chen, Jiayi Feng, Chaofeng Sha, Xin Peng, and Yiling Lou. 2024. Evaluating Large Language Models in Class-Level Code Generation. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering (Lisbon, Portugal) (ICSE ’24). Association for Computing Machinery, New York, NY, USA, Article 81, 13 pages. https://doi.org/10.1145/3597503.3639219
@@
- Feautrier (1996) Paul Feautrier. 1996. Some Efficient Solutions to the Affine Scheduling Problem Part I One-dimensional Time. International Journal of Parallel Programming 21 (Aug. 1996). https://doi.org/10.1007/BF01407835
- Feautrier (1997) Paul Feautrier. 1997. Some Efficient Solutions to the Affine Scheduling Problem Part II Multidimensional Time. International Journal of Parallel Programming 21 (Jan. 1997). https://doi.org/10.1007/BF01379404
- Github ([n. d.]) Github. [n. d.]. GitHub Copilot Your AI Pair Programmer. https://github.com/features/copilot
- Grosser et al. (2012) Tobias Grosser, Armin Groesslinger, and Christian Lengauer. 2012. Polly - Performing Polyhedral Optimizations on a Low-Level Intermediate Representation. Parallel Process. Lett. 22, 04 (Dec. 2012), 1250010. https://doi.org/10.1142/S0129626412500107
- Ivanov et al. (2024) Ivan R. Ivanov, Joachim Meyer, Aiden Grossman, William S. Moses, and Johannes Doerfert. 2024. Input-Gen: Guided Generation of Stateful Inputs for Testing, Tuning, and Training. arXiv:2406.08843 [cs.SE] https://arxiv.org/abs/2406.08843
@@
- Kind et al. (2022) Cecilia Conde Kind, Michael Canesche, and Fernando Magno Quintao Pereira. 2022. Jotai: a Methodology for the Generation of Executable C Benchmarks. Technical Report. Technical Report 02-2022. Universidade Federal de Minas Gerais.

### v2

@@ 3.2.5. Auto-Tuning
For practical reasons, ML approaches used in auto-tuning (Mustafa, 2022; Ashouri et al., 2018) mostly rely on local search methods with focused search around particular regions in the search space, after aggressively pruning the search space. Similar to the exploitation in EC algorithms, Tadashi can be setup to narrow down the search around the region for auto-tuning, based on the post-pruning range of interest.
@@ 3.2.6. ML-driven deployment of applications on new hardware
In addition to supporting the usage patterns of the ML approaches listed above, Tadashi can drive an ML approach to learnhardware specific optimisations on new hardware. After the training phase, applications ported to the new hardware can do inference on the ML model, leveraging Tadashi, to optimize the applications for the new hardware.
@@ 4.1. The App Class
The App class (Listing 2 line 3) is completely written in Python and provided mostly for convenience. It’s main functions are to provide the source_path and include_paths properties, the former being the C source file which will be transformed, and the later a list of paths used to set the required environment variable,444Namely C_INCLUDE_PATH!. which is necessary, because when PET runs a compiler pass on the input file, it will throw an error if it cannot find an included header file.
@@ 4.2. SCoPs
Accessing the Scops object (line 4) sets things in motion in the C/C++ back-end of Tadashi. The input file is passed through PET’s C source transformation function555Namely pet_transform_C_source. which detects all the SCoPs in the file, but instead of doing any transformations, the pet_scop objects are used to populate a container tadashi_scop structs. These structs store bookkeeping information and ISL objects, such as the domain, schedule and dependencies, required for the legality checks. Dependencies are computed by ISL as described in (Verdoolaege et al., 2021) and algorithm adopted from PPCG (Verdoolaege et al., 2013) is executed at this point to eliminate dead code. Tadashi also handles the management of loading multiple apps, requiring the creation, storage and deletion of multiple Scops.
@@ 4.3. Schedule Tree and Navigation
Each transformation in Tadashi is defined on a node of a schedule tree, so line 4 in Listing 2 involves the following mechanism. First, scops[i] returns a Python Scop object which stores the index of the corresponding tadashi_scop entry on the C++ side. This Scop object wraps all the C++ functions exposed to Python by calling them with a scop_idx parameter from an object method with same name. These functions include navigation operations666Namely goto_root(), goto_parent(), and goto_child(idx). to update the current node to point to its parent, its child, or to the root of the schedule tree (for the selected tadashi_scop object). The schedule_tree property on line 4 is dynamically generated, and exposes the schedule tree as a list of nodes by doing a depth-first traversal. During the traversal, we build a list of child indexes which enables Tadashi to locate any node by calling goto_child for each element in the list (implemented as the locate() method of a Node). In addition to the navigation methods, Tadashi exposes wrappers to query properties of nodes, (such as node type, number of children, partial schedule, parameters, enclosing iterators etc.), and so enables the Python Node object to emulate the ISL schedule tree and its nodes in C++.
@@ 4.4.1. Implementation Details of Individual Transformations
To merge all band node under a sequence (FULL_FUSE), we take their partial schedules, intersect them with the corresponding filters and take the union. Then we introduce a new band node on top of the sequence777Using isl_schedule_insert_partial_schedule!. and delete the original band nodes.
The 6 different SHIFT transformations, while simple on the surface, involve a tedious process involving the unpacking of the partial schedule, iterating through the sets of its domain and (re)constructing the adequate mupa888Short for isl_multi_union_pw_aff!. object.
@@ 4.5. Legality Checking
After a transformation concludes, the legality check is invoked on the same node. The legality check implements the principles described in Section 2.2. Each step of the legality check in Algorithm 1 can be implemented with a single ISL function. Algorithm 1 is applied after every transformation with the exception of SET_PARALLEL. For the legality of parallel loops the entry in corresponding to the loop should be , and this can be achieved with a slight modification to Algorithm 1: we use the partial schedule instead of the complete schedule in steps 1 and 2, and in step 4 we compute and and check if they are both empty in step 5. Figure 4 shows two examples of the process of checking legality. Checking legality starts with the dependency graph at the top, which, in case of Listing 1, is one vertex with a loop edge.
The domain and range of the map (below ), which indicates that statement instance depends on , are transformed according to the schedule .
The resulting transformed map is shown below (dotted rectangle).
If the difference between the (transformed) domain and the (transformed) range, denoted by , is lexicographicaly positive, the schedule is legal (left side, the from Figure 2(c)), otherwise the schedule is not legal (right side, the from Figure 2(d)).
@@ 4.6. Code Generation
Apart from the obvious tweaking to use the modified schedule instead of the original one, and to make it cooperate with the bookkeeping setup we use in Tadashi’s back-end, the SET_PARALLEL transformation required additional modifications to the core code generation.
@@ 5. Evaluation
We evaluate the capabilities of Tadashi by verifying functional correctnes, using two different ML solutions on top of Tadashi to showcase how Tadashi enables ML solutions to speedup different applications and benchmark suites. Then we demonstrate how Tadashi can be integrated with existing scientific computing codes, illustrating Tadashi’s potential to enabling AI-based automated code generation with guaranteed correctnesss for scientific applications, and finally measure the overhead.
5.1. Empirical Functional Testing
We provided correctness guarantees of the polyhedral method we developed for Tadashi in Section 4.5. However, we discuss in this section how we empirically do functional testing to assure the correct implementation of Tadashi. We evaluated Tadashi by applying a range of primitive (single) and composite (combined) transformations on kernels from the Polybench (Reisinger, 2024) benchmark suite.
To verify the correctness of implementation of Tadashi,
@@ 5.1. Empirical Functional Testing
first, all Polybench benchmarks were compiled and executed with the small dataset999Using -DSMALL_DATASET!. and with the output saved.101010Using -DPOLYBENCH_DUMP_ARRAYS!. In an automated manner, we generate 3 random transformations,
rollback the illegal ones, compile, run, and compare the output of the transformed program to the original.
Post-inspecting confirmed the outputs to be identical.
5.2. Using Tadashi with ML Solutions
To demonstrate the practical utility of Tadashi in real-world ML workflows, we developed and deployed three illustrative examples using our framework. The first is a simple proof-of-concept implementation of an optimal transformation search (POC-search), designed to highlight Tadashi’s ability to apply, verify and evaluate transformations: While experimenting with transformations, we noticed fission and tiling (2D and 3D) can often speed up Polybench kernels,111111Using the extra large dataset. so with a fixed tile size, we implemented a brute-force search, applying fission and tiling when ever possible, rolling back and discarding illegal results to obtain speedups presented in Figure 5. The atax, bicg, gesummv, and jacobi-1d kernels are not included since their execution time is under 0.01s and, as such, improvements on the results can be caused by noise in the CPU. A speedup of 1x indicates that no valid operations were found (cholesky, deriche, durbin, ludcmp, seidel-2d) or that the transformation did not result in a speedup (gesummv).
POC algorithm speedups
The other two are representative ML models that showcase how Tadashi enables verifiably correct code generation in the context of ML. Each of these examples illustrates different aspects of Tadashi’s design and capabilities.
5.2.1. Monte-Carlo Tree Search
We have adopted Monte Carlo Tree Search (MCTS) (Coulom, 2007) as a baseline method for discovering effective transformations and compositions of transformations. MCTS algorithm iteratively builds a search tree where nodes represent states of the code after applying specific transformations (such as tiling, fusion, or interchange), and edges represent the transformations themselves. More precisely, we break each transformation down into three steps: selecting which node to transform, which transformation to apply and which parameters (e.g., tile size) to select for a given transformation. This approach naturally handles the composition of transformations: applicability and effect of each transformation depends on previously applied ones. MCTS requires no domain-specific heuristics beyond a performance evaluation function, thus allowing it to discover non-intuitive transformations.
Furthermore we employ the Upper Confidence Bound for Trees (UCT) (Kocsis and Szepesvári, 2006) strategy to balance exploration of novel transformation sequences with exploitation of already discovered effective patterns.
Figure 6 illustrates the pace of discovery of effective transformations for a range of kernels from the Polybench suite. We excluded the loop parallelization transformation since in most cases it is easy to do manually, thus we wanted to focus on single-core performance improvement. Each chart is a single run of Tadashi on an AMD EPYC 9684X 96-Core Processor. The main cost of running Tadashi is the actual execution of the transformed code. We aimed at execution time of about 10 seconds for a given kernel, to strike a balance between reliability of measurements and reasonable turn-around time. Since the execution time can be set arbitrary, we report performance improvement over the number of evaluation steps.
Plot of speedups discovered by MCTS on different benchmarks
In line with our intuition for the POC approach, MCTS quickly discovered that as stand-alone transformations, interchange (e.g., for gemm) and tiling (e.g., for floyd-warshall) are often very effective. In terms of compositions, we see interchange or full split combined with tiling being discovered on multiple occasions. As a concrete example, transformations resulting in 6x performance improvement in trmm kernel were discovered as follows:
trs = [ [3, "FULL_SPLIT"], [1, "TILE1D", 3], [5, "INTERCHANGE"] ]
The first number is the index of a node in the polyhedral schedule tree, last number indicates the transformation parameter. More examples can be found in the examples/evaluation/data directory in the project repository.
5.2.2. Evolutionary Computation
In addition to the MCTS, we explored using a Genetic Programming (GP)-based (Vanneschi and Silva, 2023) approach to search for a good set of transformations. Unlike MCTS, GP is allowed to use all transformations provided by Tadashi. We limit the number of iterations to 20 and use a population size of 20, i.e., the algorithm does a total of 400 evaluations to obtain the final list of operations.
In this implementation, a model is represented as a list of operations to be applied to a SCoP. The models are initialized as an empty list of transformations. During training, the models can be mutated to append a randomly selected transformation to their list or lose one operation, and can also be combined with other models by picking a cutting point from both models’ lists and swapping the latter halves, creating two new models. Any invalid models generated this way are discarded.
The GP experiments are still in a preliminary stage, and while we only obtained results for two of the Polybench benchmarks (gemm and jacobi-2d), the results are promising. On gemm, GP obtained a 18.2x speedup (4.0x, without using the parallel transformation) after 12 iterations (240 evaluations). On jacobi-2d, GP obtained a 2.8x speedup after 8 iterations (160 evaluations). These results indicate that GP can beat the results previously seen in Figure 5.
5.2.3. Transformations Found by Tadashi
The following lists show two examples of the Tadashi transformations that GP found to speedup gemm and jacobi-2d by 18.2x and 2.8x, respectively. Additional details of transformations that Tadashi discover are included in the supplementary material, as permitted for accepted papers. Each row contains the node index, the transformation and its arguments.
gemm = [ [1, "FULL_SHIFT_PARAM", 2, 16], [2, "FULL_FUSE", ], [1, "SET_PARALLEL", 6], [2, "FULL_SHIFT_PARAM", 1, 48], [2, "TILE2D", 32, 32], [3, "SET_LOOP_OPT", 0, 3], ] jacobi-2d = [ [9, "TILE1D", 32], [4, "TILE2D", 32, 32], [10, "TILE2D", 32, 32], [10, "TILE1D", 32], [13, "INTERCHANGE"], ]
5.3. Integration with Scientific Applications
To evaluate the applicability of Tadashi into more realistic scientific computing scenarios, beyond simple benchmarks overspecialized for polyhedral optimization, we integrated it with two representative applications: miniAMR (Vaughan and Barrett, 2015) and FDTD electromagnetic simulation from (Sullivan, 2013). These codes, while manageable in size, reflect the complexity and structure of production-level scientific software, providing valuable insight into how Tadashi can support correctness and agency in practical settings.
miniAMR is a proxy application developed as part of the Mantevo project, designed to emulate the behavior of Adaptive Mesh Refinement (AMR) codes used in large-scale physics simulations. It features block-structured grids, dynamic refinement and coarsening, and stencil-based updates—characteristics commonly found in production AMR codes.
The FDTD electromagnetic simulation uses the Finite-Difference Time-Domain (FDTD) method. It implements a full 3D simulation kernel including absorbing boundary conditions, and serves as a self-contained, practical example of a numerical method widely used in computational electromagnetics.
The FDTD electromagnetic simulation is relatively simple and self-contained, with a clear computational kernel, which made it well suited for Tadashi-driven optimization–resulting in a 13x speedup. miniAMR, on the other hand, is significantly more complex, featuring dynamic mesh refinement and distributed data structures. Despite these challenges, Tadashi was able to produce a 1.05x (5%) speedup, demonstrating utility even in more intricate scientific codes.
The modest 1.05x speedup in miniAMR highlights the need for more sophisticated machine learning techniques and a more elaborate setup to fully exploit the potential of Tadashi in complex applications. Specifically, a more refined strategy for distributing the evaluations and measuring the runtime across multiple nodes would likely yield better scaling of ML methods. Fortunately, Tadashi fully supports such a distributed setup, leveraging tools like MPI4Py.futures to enable parallel execution and scalable optimizations across a cluster. This capability will be crucial as we continue to refine our approach and tackle even more complex scientific workloads.
5.4. Overhead
@@
The overhead measurements are done on a regular desktop with a 12 core AMD Ryzen 5 5600X 6-Core Processor and 64GiB of memory, running Arch Linux, with python 3.12.5-1, libisl 0.27-1, llvm 18.1.8-4 system packages, and Tadashi’s CMake script using the current master branch of PET121212From https://repo.or.cz/pet.git. tagged as pet 0.11.8.
@@ 6. Conclusion
While our initial experiments with machine learning methods validated the utility of Tadashi and the approach in general, they also revealed the vastness of the search space: some kernels in real applications translated into hundreds of polyhedral schedule tree nodes, each of which could be modified by a combination of numerous transformations and parameters. This realization, in turn, necessitates future work to create a solution that would allow Tadashi to run at the scale of large HPC clusters, enabling the discovery of polyhedral optimizations that go beyond those described in the literature. We leave this exploration for future work.
7. Acknowledgement
ChatGPT (OpenAI, [n. d.]) was used for Grammar and Spelling correction.
@@ Acknowledgements.
- exa ([n. d.]) [n. d.]. The Exascale Software Portfolio | Science & Technology Review. https://str.llnl.gov/past-issues/february-2021/exascale-software-portfolio
- 202 (2025) 2025. Sakana AI’s Claim That It Can ’make Things 100 Times Faster’ Has Been Verified Online and It Has Been Pointed out That It Will Actually ’Make Things 3 Times Slower’. http://gigazine.net/gsc_news/en/20250225-sakana-ai-cuda-engineer-walks-back/
@@
- Ashouri et al. (2018) Amir H. Ashouri, William Killian, John Cavazos, Gianluca Palermo, and Cristina Silvano. 2018. A Survey on Compiler Autotuning using Machine Learning. ACM Comput. Surv. 51, 5, Article 96 (sep 2018), 42 pages. doi:10.1145/3197978
@@
- Cai et al. (2023) Zonghui Cai, Xiao Yang, MengChu Zhou, Zhi-Hui Zhan, and Shangce Gao. 2023. Toward explicit control between exploration and exploitation in evolutionary algorithms: A case study of differential evolution. Information Sciences 649 (2023), 119656. doi:10.1016/j.ins.2023.119656
@@
- Coello (2022) Carlos A. Coello Coello. 2022. Constraint-handling techniques used with evolutionary algorithms. In Proceedings of the Genetic and Evolutionary Computation Conference Companion (Boston, Massachusetts) (GECCO ’22). Association for Computing Machinery, New York, NY, USA, 1310–1333. doi:10.1145/3520304.3533640
- Cotroneo et al. (2024) Domenico Cotroneo, Alessio Foggia, Cristina Improta, Pietro Liguori, and Roberto Natella. 2024. Automating the correctness assessment of AI-generated code for security contexts. Journal of Systems and Software 216 (2024), 112113. doi:10.1016/j.jss.2024.112113
- Coulom (2007) Rémi Coulom. 2007. Efficient Selectivity and Backup Operators in Monte-Carlo Tree Search. In Computers and Games, H. Jaap van den Herik, Paolo Ciancarini, and H. H. L. M. (Jeroen) Donkers (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 72–83.
- Cummins et al. (2020) Chris Cummins, Zacharias V. Fisches, Tal Ben-Nun, Torsten Hoefler, and Hugh Leather. 2020. ProGraML: Graph-based Deep Learning for Program Optimization and Analysis. (March 2020). doi:10.48550/arXiv.2003.10536
@@
- Cummins et al. (2024) Chris Cummins, Volker Seeker, Dejan Grubisic, Baptiste Roziere, Jonas Gehring, Gabriel Synnaeve, and Hugh Leather. 2024. Meta Large Language Model Compiler: Foundation Models of Compiler Optimization. doi:10.48550/arXiv.2407.02524 arXiv:2407.02524 [cs]
- Dehaerne et al. (2022) Enrique Dehaerne, Bappaditya Dey, Sandip Halder, Stefan De Gendt, and Wannes Meert. 2022. Code Generation Using Machine Learning: A Systematic Review. IEEE Access 10 (2022), 82434–82455. doi:10.1109/ACCESS.2022.3196347
- Du et al. (2024) Xueying Du, Mingwei Liu, Kaixin Wang, Hanlin Wang, Junwei Liu, Yixuan Chen, Jiayi Feng, Chaofeng Sha, Xin Peng, and Yiling Lou. 2024. Evaluating Large Language Models in Class-Level Code Generation. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering (Lisbon, Portugal) (ICSE ’24). Association for Computing Machinery, New York, NY, USA, Article 81, 13 pages. doi:10.1145/3597503.3639219
@@
- Feautrier (1996) Paul Feautrier. 1996. Some Efficient Solutions to the Affine Scheduling Problem Part I One-dimensional Time. International Journal of Parallel Programming 21 (Aug. 1996). doi:10.1007/BF01407835
- Feautrier (1997) Paul Feautrier. 1997. Some Efficient Solutions to the Affine Scheduling Problem Part II Multidimensional Time. International Journal of Parallel Programming 21 (Jan. 1997). doi:10.1007/BF01379404
- Github (2022) Github. 2022. GitHub Copilot Your AI Pair Programmer. https://github.com/features/copilot
- Grosser et al. (2012) Tobias Grosser, Armin Groesslinger, and Christian Lengauer. 2012. Polly - Performing Polyhedral Optimizations on a Low-Level Intermediate Representation. Parallel Process. Lett. 22, 04 (Dec. 2012), 1250010. doi:10.1142/S0129626412500107
- Heroux (2024) Michael A. Heroux. 2024. Scalable Delivery of Scalable Libraries and Tools: How ECP Delivered a Software Ecosystem for Exascale and Beyond. Computing in Science & Engineering 26, 1 (Jan. 2024), 9–18. doi:10.1109/MCSE.2024.3384937
@@
- Kocsis and Szepesvári (2006) Levente Kocsis and Csaba Szepesvári. 2006. Bandit based monte-carlo planning. In Proceedings of the 17th European Conference on Machine Learning (Berlin, Germany) (ECML’06). Springer-Verlag, Berlin, Heidelberg, 282–293. doi:10.1007/11871842_29